Find the value of k so that the lines through the given points are parallel. Line 1:(-1,-9)and(-6,-6) Line 2:(-7,K)and(0,-2)

Answer by mananth(16946)

(-1,-9)and(-6,-6) Line 2:(-7,K)and(0,-2)
slope of line 1 = slope of line 2
m= (y2-y1)/(x2-x1)
-6-(-9)/-6-(-1)
=3/-5
...
(-7,K)and(0,-2)
m=-2-k/0-(-7)
m=-2-k/7
..
3/-5 =(-2-k)/7
-5(-2-k)=7*3
10+k=21
k=21-10
k=11
...
m.ananth@hotmail.ca